A non interest-bearing asset has a spot price of USD 100. It costs 1% per annum (on a simple rate basis) to store the asset. Simple interest rates are 5% per annum. What is the fair one-year forward price?
Answer #1

fair one-year forward price = spot price*e(r+q)*T

r = interest rate; q = storage cost; T = time period

fair one-year forward price = $100*e(0.05+0.01)*1 = $100*e0.06 = $100*1.0618 = $106.18